An analysis of equity of Hahn Corporation as of January 1, 2020, is as follows: Share capital—ordinary, par value P20; authorized 100,000 shares; issued and outstanding 90,000 shares P1,800,000 Share premium—ordinary 900,000 Retained earnings 760,000 Total P3,460,000 Hahn uses the cost method of accounting for treasury shares and during 2020 entered into the following transactions: Acquired 2,500 of its shares for P75,000. Sold 2,000 treasury shares at P35 per share. Sold the remaining treasury shares at P20 per share. Assuming no other equity transactions occurred during 2020, what should Hahn report at December 31, 2020, as total share premium?
Answer #1
Working Notes:
Calculation of Purchase Price of Treasury Stock
Purchase value of Treasury Stock 2,500 Shares =                75,000
Purchase Price per Shares = P 75,000 / 2,500 =                        30
Solution:
Calculation of Total Share Premium as on December 31. 2020
Beginning Balance of Share Premium - Ordinary              9,00,000
Shares Premium on Treasury Stock
2000 Shares are sold at Premium of P 5 (35 - 30) = (2000 X 5)                10,000
Remainning 500 Shares at Loss of P 10 (30 -20 ) = (500 X 10)                 -5,000
Share Premium - Treasury Stock                    5,000
Total Share Premium              9,05,000
Answer = Hahn Reported at December 31, 2020, a total Share Premoum = P 905,000